OpenStudy (anonymous):

The price of an all-access ticket to the fair is $42.99 before tax. Roland bought the all-access ticket for 40% off. He then paid 7% sales tax on the discounted price. Part A: How much did Roland pay in sales tax? Show all work and steps in your solution. Part B: What is the total amount that Roland paid for the ticket? Show all work and steps in your solution

OpenStudy (anonymous):

@jim_thompson5910 my sale tax is 1.74

OpenStudy (anonymous):

40%*42.99-17.19 =24.99 then I find t

OpenStudy (anonymous):

7% of 24.99

jimthompson5910 (jim_thompson5910):

x = original price of ticket 40% off means he pays 60% of the full price, so 0.60x 7% tax means he has to add an additional 0.07*(0.60x) = 0.042x to the final cost so the total cost is 0.60x+0.042x = 0.642x

jimthompson5910 (jim_thompson5910):

hopefully you're seeing how I'm getting all this?

OpenStudy (anonymous):

yes I do

jimthompson5910 (jim_thompson5910):

"The price of an all-access ticket to the fair is $42.99 before tax" so ignore the tax part and just focus on 0.60x

jimthompson5910 (jim_thompson5910):

0.60x = 42.99 x = ???

OpenStudy (anonymous):

I was thinking about the reduce

jimthompson5910 (jim_thompson5910):

hmm I'm not sure if $42.99 is the price BEFORE the 40% reduction? or AFTER the 40% reduction?

OpenStudy (anonymous):

that was confusing

jimthompson5910 (jim_thompson5910):

yeah I'm stuck too

OpenStudy (anonymous):

guess tax is always paid last

jimthompson5910 (jim_thompson5910):

yeah true

jimthompson5910 (jim_thompson5910):

but I'm still not sure if $42.99 is the original cost or the reduced cost

OpenStudy (anonymous):

after re-reading i think its the original and just to show how they arrive at it base on the calculation

jimthompson5910 (jim_thompson5910):

ok so if $42.99 is the original price before the 40% reduction then 60% of 42.99 = 0.60*42.99 = ??

OpenStudy (anonymous):

71.65

jimthompson5910 (jim_thompson5910):

.6*42.99 = 25.794

OpenStudy (anonymous):

25.79

jimthompson5910 (jim_thompson5910):

original price before reduction = $42.99 price after reduction = $25.79

jimthompson5910 (jim_thompson5910):

now take 7% of that reduced price to find how much tax is paid

OpenStudy (anonymous):

1.81

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

23.98 he paid for tickets

jimthompson5910 (jim_thompson5910):

total price = (price after reduction) + (tax)

OpenStudy (anonymous):

27.60

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks for your help

jimthompson5910 (jim_thompson5910):

no problem